# Turning Trash into Treasure: Accelerating Inference of Large Language Models with Token Recycling

### v1

@@ Abstract
The rapid growth in the parameters of large language models (LLMs) has made inference latency a fundamental bottleneck, limiting broader application of LLMs. Speculative decoding represents a lossless approach to accelerate inference through a guess-and-verify paradigm, leveraging the parallel capabilities of modern hardware. Some speculative decoding methods rely on additional structures to guess draft tokens, such as small models or parameter-efficient architectures , which need extra training before use. Alternatively, retrieval-based train-free techniques build libraries from pre-existing corpora or by n-gram generation. However, they face challenges like large storage requirements, time-consuming retrieval, and limited adaptability. Observing that candidate tokens generated during the decoding process are likely to reoccur in future sequences, we propose Token Recycling. This approach stores candidate tokens in an adjacency matrix and employs a breadth-first search (BFS)-like algorithm on the matrix to construct a draft tree. The tree is then validated through tree attention. New candidate tokens from the decoding process are then used to update the matrix. Token Recycling requires <2MB of additional storage and achieves approximately 2x speedup across all sizes of LLMs. It significantly outperforms existing train-free methods by 30% and even a training method by 25%. It can be directly applied to any existing LLMs and tasks without the need for adaptation.
Figure 1: A comparison of typical speculative decoding and Token Recycling. Typical speculative decoding drafts some tokens and verifies them in parallel in one decoding step. Unlike speculative decoding which discards candidate tokens, Token Recycling stores them in an adjacency matrix. In future generations, draft tokens are retrieved from the matrix which is updated with new candidate tokens. Token Recycling effectively recycles tokens in the decoding process.
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al. 2020; Gemini Team et al. 2023; Touvron et al. 2023; Meta 2024) have made significant contributions to the development of artificial intelligence, becoming the foundation of numerous applications such as chatbots, code assistants, agents (OpenAI 2023; Chen et al. 2021; Wang et al. 2024a). However, due to the auto-regressive decoding strategy, LLMs can only generate a single token at each decoding step, leading to slow response times for queries. The latency mainly comes from transferring billions of parameters from High-Bandwidth Memory to the accelerator’s cache at each decoding step, rather than arithmetic computations (Kim et al. 2024; Shazeer 2019; Cai et al. 2024).
@@
Many approaches (Xu et al. 2024; Frantar and Alistarh 2023; Dao 2024; DeepSeek-AI 2024) seek to reduce the latency, with speculative decoding as a key lossless technique. This approach employ a guess and verify process to obtain multiple tokens during a single decoding step  (Chen et al. 2023; Leviathan, Kalman, and Matias 2023; Miao et al. 2024; Xia et al. 2023). It first speculates several subsequent draft tokens and then verifies them using the original LLMs. The time cost of verification on multiple tokens is comparable to that of generating one token due to the high parallel performance of accelerators. Once some draft tokens are correct, the whole decoding steps can be greatly reduced while maintaining the quality. To fully utilize the parallel capability of accelerators, tree attention slightly adjust the attention mask to verify multiple draft tokens continuations in one model forward (Cai et al. 2024; Miao et al. 2024).
@@
Speculative decoding aims not only to maintain quality but also to minimize the cost of speculation. Some researchers construct additional model structures to guess the draft tokens, including small draft models (Leviathan, Kalman, and Matias 2023; Chen et al. 2023) and parameter-efficient structures (Cai et al. 2024; Lin et al. 2024). However, these approaches require resources for additional training on each LLM. The typical approach to achieve train-free speculative decoding is retrieve-based. In this case, a retrieval library is pre-defined to obtain tokens following the suffix of current content as draft tokens. Several methods have been proposed in this category, each with its trade-offs: (i) REST (He et al. 2023) transforms existing corpora into a retrieval library, but the storage is large, retrieval is time-consuming, and the library lacks flexibility as it’s static to any queries. (ii) PLD (Saxena 2023) only retrieves the previous content with minimal cost. However, it can not predict new tokens or new combinations of tokens. (iii) Lookhead (Fu et al. 2024) construct and update an n-gram library through LLMs decoding n times. However, LLMs have to generate n-grams while in inference, causing low efficiency.
@@
Furthermore, all these speculative decoding approaches fail to fully utilize candidate tokens, which are multiple possible next tokens generated by LLMs at each decoding step. Typically, only the top-1 candidate token from accepted tokens is selected as the output, while other candidate tokens, including all candidate tokens from rejected tokens, are discarded, such as ‘k’ and ‘keys’ in Figure 1. However, we observe that when current input tokens reappear in future generations, the following tokens could be candidate tokens generated several steps prior. Based on the observation, we propose Token Recycling, a method which utilizes candidate tokens as draft tokens. It stores candidate tokens in an adjacency matrix. Before each decoding step, a BFS-like approach retrieves a draft tree from the matrix, which is then verified using tree attention. Once verified, the newly generated candidate tokens update the matrix. (i) The matrix provides a flexible retrieval library that is tailored to each query and offers low retrieval costs due to its small size (<2MB). (ii) Compared to using the previous content solely, candidate tokens naturally include more tokens, providing many possible continuations. (iii) The construction and update of our library (matrix) utilize the ‘trash’ tokens without requiring any additional generation.
@@
We conduct comprehensive experiments on general benchmark SpecBench (Xia et al. 2024), and specialized dataset on code domain, MBPP  (Chen et al. 2021) with different sizes of Vicuna (Zheng et al. 2023) and Code Llama (Roziere et al. 2023) (7b, 13b, 33b/34b). The results show that Token Recycling greatly exceeds previous train-free approaches, and improves more than 31% on all sizes. The speed-up ratio even exceeds the widely used training approach–Medusa, demonstrating its high efficiency.
@@
  • We observe that candidate tokens can be reused as draft tokens in subsequent sequences. Based on this observation, we propose a train-free speculative decoding method, Token Recycling, which saves candidate tokens and retrieves draft tokens based on the adjacency matrix.
  • The matrix requires minimal storage space (<2MB) with a low retrieval cost, and covers many new tokens. Continuously updating provides a dynamic retrieval space, but doesn’t need additional generation.
  • The experimental results show that Token Recycling achieves approximately 2x speedup on all sizes of LLMs. It improves more than 31% compared to previous train-free approaches and achieves a new SOTA, even exceeding a training approach.
@@ 2 Background
In this section, we overview the speculative decoding. We begin with a formal definition of auto-regressive decoding. Next, we explore speculative decoding, highlighting the guess-and-verify strategy. Finally, we introduce tree attention, which verifies multiple draft sequences simultaneously.
@@ 2.1 Auto-Regressive Decoding
Auto-Regressive (AR) is the default decoding strategy of LLMs. At each decoding step $t$, LLMs calculate the probability distribution of the next token given the current content $s=(x_{0},x_{1},\cdots,x_{t})$ which $x_{i}\in\mathcal{V}$:
@@
Here $\mathcal{V}$ represents the vocabulary and $\theta$ represents the parameters of the LLM. The next token is selected from $p_{t+1}$ depending on the sampling methods. Followed Kou et al. (2024), we focus on greedy decoding in this paper. The next token is as follows:
@@
Candidate tokens refer to the top-$k$ tokens with the highest probabilities
@@
where $k$ is the number of candidate tokens. Here, $\text{argtop}k(\cdot)$ is a custom operator that returns the indices of the top-$k$ highest values in the probability distribution $p_{t+1}$.
Figure 2: An overview of Token Recycling. The adjacency matrix, initialized by inheriting from the previous query, stores candidate tokens. Token Recycling first retrieves a draft tree from the matrix based on the last token of the current content. The tree is then compressed into a merged sequence with a corresponding tree attention mask and sent to the LLM for a forward pass. After processing, all possible draft sequences are extracted and verified. The longest correct sequence is selected and added to the content. Finally, the top-k candidate tokens are used to update the matrix for the next iteration.
@@ 2.2 Speculative Decoding
Through the guess and verify approach, speculative decoding effectively utilizes the parallel capability of GPUs to accelerate inference. Given $s$, it first guesses $n$ subsequent draft tokens $(\tilde{x}_{t+1},\cdots,\tilde{x}_{t+n})$. The combination $(s,\tilde{x}_{t+1},\cdots,\tilde{x}_{t+n})$ is then sent to LLMs for one forward pass, resulting in:
@@
$\tilde{p}_{t+i}$ $=P(x\mid s,\tilde{x}_{t+1},\ldots,\tilde{x}_{t+i-1};\theta),\quad\text{for }i=2,\ldots,n.$
$p_{t+1}$ is the same as AR decoding so the ground-truth $x_{t+1}$ can be selected using ‘argmax’. If the draft token $\tilde{x}_{t+1}$ matches $x_{t+1}$, then $\tilde{p}_{t+2}$ is assumed to identical to $p_{t+2}$ in AR decoding. Thus, the next token is selected by: $x_{t+2}=\text{argmax}\ \tilde{p}_{t+2}$. This verify process continues until the draft token does not match the selected token, indicated by:
@@
Ultimately, $j$ new tokens are confirmed in one forward pass. The time cost of one forward pass with $(s,\tilde{x}_{t+1},\cdots,\tilde{x}_{t+n})$ is nearly the same as with $s$ due to the high parallel performance of accelerators. Figure 1 shows an example. Consider the current content ending with the token ‘for’ and the draft tokens are [‘i’, ‘in’, ‘range’, ‘(’]. After the forward pass, the output token for ‘for’ is ‘i’, and for ‘i’ is ‘in’, but for ‘in’ is ‘zip’, which fails to match the guessed ‘range’. ‘i’ and ‘in’ are accepted tokens while the others are rejected. Despite ‘zip’ not matching ‘range’, it is still a valid output generated by the LLM based on the correct content, resulting in three tokens [‘i’, ‘in’, ‘zip’] being confirmed in one forward pass. Overall, speculative decoding effectively accelerates inference while maintaining the same output as AR decoding.
2.3 Tree Attention
Traditional causal attention masks are designed for linear sequences, where each token attends to all previous tokens. This design restricts speculative decoding to verifying one sequence at a time. However, as the sequence lengthens during draft token generation, the number of potential continuations increases. For example, in the draft tree in Figure 2, the token following ‘guest’ could be ‘speaker’ or ‘speak’, while both ‘at’ and ‘for’ could follow ‘speaker’. This creates a need to verify multiple draft sequences simultaneously.
Tree attention modifies the attention mask to address this. It compresses multiple sequences into a single merged sequence, such as [‘guest’, ‘speaker’, ‘speak’, ‘at’, ‘for’, ‘ings’], while preserving a tree structure in the tree attention mask. Each child node attends only to its parent nodes, preventing sibling tokens from interfering with each other. After the LLM processes the merged sequence, all possible sequences such as ‘guest speaker’, ‘guest speaker at’, ‘guest speaker for’, and ‘guest speak’ and so on, along with their corresponding output tokens are extracted based on the tree structure and verified in parallel. The longest correct sequence is selected as the final output. Tree attention enables speculative decoding to verify multiple draft sequences in one decoding step, significantly enhancing efficiency while maintaining output quality.
In summary, speculative decoding, through guess and verify and tree attention, improves the inference latency robustly and efficiently compared to AR decoding.
@@ 3 Methodology
Figure 2 provides an overview of Token Recycling. It leverages a hot-start adjacency matrix to store candidate tokens and employs a BFS-like algorithm to construct a draft tree. It utilizes tree attention to verify draft sequences and continuously updates the matrix with new candidate tokens generated during the decoding process.
@@ 3.1 Adjacency Matrix Initialization
The adjacency matrix $\mathcal{M}$ is a key component in Token Recycling, used to store top-$k$ candidate tokens for each token in the vocabulary:
@@
Initially, all elements are set to zero, meaning that a token must appear in draft tokens before it has valid candidate tokens. This initialization leads to the matrix starting with limited predictive capability, potentially causing inefficiencies during the early stages of inference. To mitigate this limitation, we implement a hot start strategy. This involves inheriting the matrix from the previous query, thereby leveraging existing knowledge. Even if queries differ in the domain, candidate tokens often include common expressions and patterns that frequently appear across various queries. Nevertheless, as the first query lacks a preceding query, the matrix is set to zero. Consequently, hot start ensures that the matrix has a broader starting point, covering a wide range of potential continuations.
@@ 3.2 Draft Tree Retrieval
The adjacency matrix $\mathcal{M}$ stores candidate tokens, which can be used as draft tokens when their corresponding tokens appear later. Directly using the matrix could only determine the immediate next token, such as finding ‘speaker’ following ‘guest’ (see Figure 2). Even if ‘speaker’ is correct, it only slightly improves upon AR decoding, adding just one additional token. In fact, the matrix also holds possible continuations for these candidate tokens, suggesting subsequent tokens like ‘at’ following ‘speaker’. Extending the sequence step by step allows for longer draft sequences. Furthermore, by storing top-$k$ candidate tokens, multiple potential continuations can be explored in parallel for each token, such as ‘at’, and ‘for’ following ‘speaker’. This breadth-first search (BFS)-like process enables the exploration of the sequence length and diverse continuations while constructing a draft tree with the adjacency matrix. Inspired by Cai et al. (2024), the tree is static and imbalanced. The complete tree structure is shown in Appendix 8.1.
  • Static: The number of children for each node remains constant across all decoding steps. This fixed structure facilitates pre-processing, such as preparing attention masks, and enables efficient parallel operations during layer traversal. By avoiding the need to individually traverse each node, it significantly reduces retrieval time.
  • Imbalance: Nodes positioned earlier in each layer have more children and extend deeper. Since candidate tokens are ordered by their probabilities in the matrix, this design prioritizes the most probable continuations, allocating computational resources to higher confident paths.
The BFS-like approach for retrieving the draft tree begins with the matrix $\mathcal{M}$ and the tree structure $Tree$. The retrieval process starts from the root, which is the last token of current content. Take the token ‘guest’ in Figure 2 as an example. As the root forms the first layer, all candidate tokens for ‘guest’ are extracted from $\mathcal{M}$, resulting in [‘speaker’, ‘speak’, ‘Spe’]. According to $Tree$, the first layer allows each token to have two children, Therefore, ‘speaker’ and ‘speak’, which have the top-2 probabilities, are selected and added to the second layer. The process then proceeds to expand a new layer starting from [‘speaker’] and [‘speak’]. All their candidate tokens are retrieved in parallel, resulting in [‘at’, ‘for’, ‘is’] and [‘ings’, ‘in’, ‘ers’]. $Tree$ specifies that the first node in this layer (‘speaker’) can have two children, while the subsequent node (‘speak’) can only have one child. Consequently, the new layer tokens are [‘at’, ‘for’], and [‘ings’]. This process repeats until the specified depth is reached. Finally, a merged sequence $S$ is constructed through traversing the draft tree by layers. The detailed Algorithm 1 is provided in Appendix 8.1.
This retrieval method achieves the desired length and variety of the draft sequences, resulting in a single merged sequence that can be sent to the LLM.

### v2

# Xuanyu Zhang2,Qing Yang2, Dongliang Xu2

@@ Abstract
Massive parameters of LLMs have made inference latency a fundamental bottleneck. Speculative decoding represents a lossless approach to accelerate inference through a guess-and-verify paradigm. Some methods rely on additional architectures to guess draft tokens, which need extra training before use. Alternatively, retrieval-based training-free techniques build libraries from pre-existing corpora or by n-gram generation. However, they face challenges like large storage requirements, time-consuming retrieval, and limited adaptability. Observing that candidate tokens generated during the decoding process are likely to reoccur in future sequences, we propose Token Recycling. It stores candidate tokens in an adjacency matrix and employs a breadth-first-search (BFS)-like algorithm to construct a draft tree, which is then validated through tree attention. New candidate tokens from the decoding process are then used to update the matrix. Token Recycling requires <2MB of additional storage and achieves approximately 2x speedup across all sizes of LLMs. It significantly outperforms existing train-free methods by 30% and even a widely recognized training method by 25%.
[Uncaptioned image] Turning Trash into Treasure: Accelerating Inference of Large Language Models with Token Recycling
Xianzhen Luo1, Yixuan Wang1, Qingfu Zhu1thanks: Corresponding author, Zhiming Zhang1, Xuanyu Zhang2,Qing Yang2, Dongliang Xu2 1Harbin Institute of Technology, Harbin, China 2Du Xiaoman (Beijing) Science Technology Co., Ltd. {xzluo, wyx, qfzhu, zmzhang}@ir.hit.edu.cn {zhangxuanyu, yangqing, xudongliang}@duxiaoman.com
Figure 1: A comparison of typical speculative decoding and Token Recycling (TR). Typical methods draft some tokens and verify them in parallel in one decoding step. Unlike other methods that discard candidate tokens, TR stores them in an adjacency matrix. In future generations, draft tokens are retrieved from the matrix which is updated with new candidate tokens. TR effectively recycles tokens in the decoding process.
@@ 1 Introduction
Large Language Models (LLMs) Brown et al. (2020); Gemini Team et al. (2023); Touvron et al. (2023); Meta (2024) have becoming the foundation of numerous applications such as chatbots, code assistants, and agents OpenAI (2023); Chen et al. (2021); Wang et al. (2024a). However, due to the auto-regressive decoding strategy, LLMs can only generate a single token at each decoding step, leading to high inference latency Brown et al. (2020). The latency mainly comes from transferring billions of parameters from high bandwidth memory to the accelerator cache at each decoding step, rather than arithmetic computations Kim et al. (2024); Shazeer (2019); Cai et al. (2024).
@@
Many approaches Xu et al. (2024); Frantar and Alistarh (2023); Dao (2024); DeepSeek-AI (2024) seek to reduce the latency, with speculative decoding as a key lossless technique. This approach employs a guess and verify process to obtain multiple tokens during a single decoding step Chen et al. (2023); Leviathan et al. (2023); Miao et al. (2024); Xia et al. (2023). It first speculates several subsequent draft tokens and then verifies them using the original LLMs. The time cost of verification on multiple tokens is comparable to that of generating one token due to the high parallelism of accelerators. Once some draft tokens are correct, the decoding steps is significantly shortened without sacrificing quality. To fully utilize the parallelism of accelerators, tree attention slightly adjust the attention mask to verify multiple token sequences in one model forward Cai et al. (2024); Miao et al. (2024).
@@
For effective acceleration, speculative decoding must ensure accurate draft predictions while keeping speculation overhead low. Additional model architectures are constructed to guess the draft tokens, including small draft models Leviathan et al. (2023); Chen et al. (2023) and parameter-efficient structures Cai et al. (2024); Lin et al. (2024). However, these approaches require resources for additional training on each LLM. The typical approach to achieve train-free speculative decoding is retrieve-based. In this case, a retrieval library is pre-defined to obtain tokens following the suffix of current content as draft tokens. Several methods have been proposed in this category, each with its trade-offs: (i) REST He et al. (2023) transforms existing corpora into a retrieval library, but the storage is large, retrieval is time-consuming, and the library lacks flexibility as it’s static to any queries. (ii) PLD Saxena (2023) only retrieves the previous content with minimal cost. However, it can not predict new tokens or new token combinations. (iii) Lookhead Fu et al. (2024) construct and update an n-gram library by decoding n times with LLMs. However, LLMs have to generate n-grams while in inference, causing low efficiency.
@@
Furthermore, all speculative decoding approaches fail to fully utilize candidate tokens, which are multiple possible next tokens generated by LLMs at each decoding step. In greedy decoding, only the top-1 candidate token of accepted tokens is selected as the output, while other candidate tokens, including all candidate tokens from rejected tokens, are discarded, such as ‘k’ and ‘keys’ in Figure 1. However, we observe that when current input tokens reappear in future generations, the following tokens could be candidate tokens generated several steps prior. Based on the observation, we propose Token Recycling (TR), which utilizes candidate tokens as draft tokens. It stores candidate tokens in an adjacency matrix. Before each decoding step, a BFS-like approach retrieves a draft tree from the matrix, which is then verified using tree attention. Once verified, the newly generated candidate tokens update the matrix. (i) The matrix provides a flexible retrieval library that is tailored to each query and offers low retrieval costs due to its small size (<2MB). (ii) Compared to using the previous content solely, candidate tokens naturally include more tokens, providing many possible continuations. (iii) The construction and update of our library (matrix) utilize the ‘trash’ tokens without requiring any additional generation.
@@
We conduct comprehensive experiments on general benchmark SpecBench Xia et al. (2024), and specialized dataset on code domain, MBPP  Chen et al. (2021) with Vicuna Zheng et al. (2023) and Code Llama Roziere et al. (2023) . The results show that TR greatly exceeds previous train-free approaches, and improves more than 30% on all sizes (7b, 13b, 33b/34b). The speed-up ratio even exceeds the widely used training approach–Medusa, demonstrating its high efficiency.
@@
  • A plug-and-play speculative decoding method, Token Recycling is proposed. It firstly recognizes the value of ‘trash’ tokens and converts them into ‘treasure’ tokens for acceleration.
  • TR requires minimal storage space (<2MB) with a low retrieval cost and covers many new tokens. Continuously updating provides a dynamic retrieval space.
  • TR achieves approximately 2x speedup on all sizes of LLMs. It achieves a new SOTA with an improvement greater than 31% compared to previous train-free approaches and even exceeding a training approach.
@@ 2 Background
In this section, we overview the speculative decoding. We first define auto-regressive (AR) decoding formally, then discuss speculative decoding, focusing on two key strategies: guess-and-verify and tree attention.
@@ 2.1 Auto-Regressive Decoding
AR is the default decoding strategy of LLMs. At each step $t$, LLMs calculate the probability distribution of the next token given the current content $s=(x_{0},x_{1},\cdots,x_{t})$ which $x_{i}\in\mathcal{V}$:
@@
Here $\mathcal{V}$ is the vocabulary and $\theta$ denotes LLM parameters. The next token is selected from $p_{t+1}$ based on the sampling method. Followed Kou et al. (2024), we focus on greedy decoding in this paper, where the next token is:
@@
Candidate tokens are the top-$k$ tokens with the highest probabilities
@@
where $k$ is the number of candidate tokens, $\text{argtop}k(\cdot)$ returns the indices of the top-$k$ highest values in $p_{t+1}$ and $x_{t+1}^{0}=x_{t+1}$.
@@ 2.2 Speculative Decoding
Guess and Verify
Speculative decoding effectively utilizes the parallel capability of accelerators. Given $s$, it first guesses $n$ subsequent draft tokens $(\tilde{x}_{t+1},\cdots,\tilde{x}_{t+n})$. The combination $(s,\tilde{x}_{t+1},\cdots,\tilde{x}_{t+n})$ is then sent to LLMs for one forward pass, resulting in:
@@
$\tilde{p}_{t+i}$ $=P(x\mid s,\tilde{x}_{t+1},\ldots,\tilde{x}_{t+i-1};\theta),i=2,\ldots,n.$
$p_{t+1}$ is the same as AR decoding so the ground truth $x_{t+1}$ is determinable. If the draft token $\tilde{x}_{t+1}$ matches $x_{t+1}$, then $\tilde{p}_{t+2}$ is assumed to identical to $p_{t+2}$. Thus, the next ground truth is selected: $x_{t+2}=\text{argmax}\ \tilde{p}_{t+2}$. This verification process continues until the draft token does not match the ground truth, indicated by:
@@
Ultimately, $j$ new tokens are confirmed in one forward pass. The time cost of one forward pass with $(s,\tilde{x}_{t+1},\cdots,\tilde{x}_{t+n})$ is nearly the same as with $s$ due to the high parallel performance of accelerators. Figure 1 shows an example. The draft tokens are [‘i’, ‘in’, ‘range’, ‘(’] and the output tokens are [‘i’, ‘in’, ‘zip’, ‘(’, ‘xs’] after the forward pass. Though ‘zip’ fails to match ‘range’, three tokens [‘i’, ‘in’, ‘zip’] are confirmed in one forward pass.
Figure 2: An overview of Token Recycling (TR). The adjacency matrix, initialized by the existing matrix, stores candidate tokens. TR first retrieves a draft tree from the matrix which is then verified through tree attention. After add the longest correct sequence to the content, the new top-k candidate tokens update the matrix.
Tree Attention
Traditional causal attention masks are designed for linear sequences, which restricts speculative decoding to verifying one sequence at a time. However, as the sequence lengthens during draft token generation, the number of potential continuations increases. For example, in the draft tree in Figure 2, the token following ‘guest’ could be ‘speaker’ or ‘speak’. Tree attention modifies the attention mask to verify multiple draft sequences simultaneously. It compresses multiple sequences into a single merged sequence, such as [‘guest’, ‘speaker’, ‘speak’], while preserving the tree structure through tree attention mask. Each child node attends only to its parent nodes, preventing sibling tokens from interfering with each other. After the LLM processes the merged sequence, all possible sequences such as ‘guest speaker’ and ‘guest speak’, along with their corresponding output tokens are extracted based on the tree structure and verified in parallel. The longest correct sequence is selected as the final output. In rare cases, when tokens have identical probabilities, tree attention and AR decoding may select different tokens, but this affects the response quality minimally. The detailed explanation is in Appendix A.1.
In summary, speculative decoding, through guess and verify and tree attention, improves the inference latency robustly and efficiently.
@@ 3 Methodology
Figure 2 provides an overview of Token Recycling (TR). It leverages a hot-start adjacency matrix to store candidate tokens and employs a BFS-like algorithm to construct a draft tree. It utilizes tree attention to verify draft sequences and continuously updates the matrix with new candidate tokens generated during the decoding process.
@@ 3.1 Adjacency Matrix Initialization
The adjacency matrix $\mathcal{M}$ is a key component in TR, used to store top-$k$ candidate tokens for each token in the vocabulary:
@@
Initially, all elements are set to zero, meaning that a token must appear in draft tokens before it has valid candidate tokens. This initialization leads to the matrix starting with limited predictive capability, potentially causing inefficiencies during the early stages of inference. To mitigate this limitation, we implement a hot start strategy. This involves continuing to use the existing matrix, thereby leveraging prior knowledge. Even if queries differ in the domain, candidate tokens often include common expressions and patterns that frequently appear across various queries. Consequently, hot start ensures that the matrix has a broader starting point, covering a wide range of potential continuations.
@@ 3.2 Draft Tree Retrieval
The adjacency matrix $\mathcal{M}$ stores candidate tokens, which can be used as draft tokens when their corresponding tokens appear later. Directly using the matrix could only determine the immediate next token, such as finding ‘speaker’ following ‘guest’ (see Figure 2). Even if ‘speaker’ is correct, it only slightly improves upon AR decoding, adding just one additional token. In fact, the matrix also holds possible continuations for these candidate tokens, suggesting subsequent tokens like ‘at’ following ‘speaker’. Extending the sequence step by step allows for longer draft sequences. Furthermore, by storing top-$k$ candidate tokens, multiple potential continuations can be explored in parallel for each token, such as ‘at’ and ‘for’ following ‘speaker’. This BFS process enables the construction of a draft tree with only the adjacency matrix, which can be directly applied to tree attention.
Unlike a complete BFS, we use heuristic rules to define a static and imbalanced tree structure. This tree structure and its construction process are detailed in the Appendix 1. Static: The number of children for each node remains constant across all decoding steps, which facilitates pre-processing and enables efficient parallel operations during layer traversal. Avoiding the need to traverse each node individually significantly reduces retrieval time. Imbalance: Nodes positioned earlier in each layer have more children and extend deeper. This allocates computational resources to the most probable continuations since candidate tokens are ordered by probabilities in the matrix.
The BFS-like approach for retrieving the draft tree begins with the matrix $\mathcal{M}$ and the tree structure $Tree$. The root is the last token of current content, like ‘guest’ in Figure 2. As the root forms the first layer, all candidate tokens for ‘guest’ are extracted from $\mathcal{M}$, resulting in [‘speaker’, ‘speak’, ‘Spe’]. According to $Tree$, the first layer allows each token to have two children, Therefore, ‘speaker’ and ‘speak’, which have the top-2 probabilities, are added to the second layer. The process then proceeds to expand a new layer. All candidate tokens of the second layer are retrieved in parallel, resulting in [‘at’, ‘for’, ‘is’] and [‘ings’, ‘in’, ‘ers’]. $Tree$ specifies that the first node (‘speaker’) can have two children, while the subsequent node (‘speak’) can only have one child. Consequently, the new layer tokens are [‘at’, ‘for’], and [‘ings’]. This process repeats until the specified depth is reached. The detailed Algorithm 1 is provided in Appendix 1.
This retrieval method constructs a draft tree effectively and efficiently with the desired length and variety, which can later be verified by tree attention.